OpenStudy (anonymous):

Will Fan and give medal for help? ^^ Solve, using linear combination. 3x – y = 4 2x + y = 6 Answers: (–1, –7) (0, –4) (2, 2) (3, 0)

OpenStudy (anonymous):

Rearrange both equations to get y on its own, then you can set them equal to each other. Ignore the y parts for now and rearrange to get x. After that you can put x back in to the equations for y (check both give the same answer).

OpenStudy (anonymous):

huh? :o @broken_symmetry

OpenStudy (anonymous):

Ok, do the first part. Rearrange your two equations so they read y = _____

OpenStudy (ranga):

Just add the two equations. That will get rid of y and you can solve for x. Then pt x in first equation and solve for y.

OpenStudy (ranga):

3x – y = 4 2x + y = 6 Add the two equations: 5x = 10 divide both sides by 5 x = 10/5 = ?

OpenStudy (anonymous):

If you have a matrix calculator, plug the numbers in and do matrix A^-1*matrix B.

OpenStudy (the_fizicx99):

-y +y cancel out its alot easier, so its just 5x = 10 5x = 10 /5 /5 x = ?